### v2

# Design and Development of Portable RPC-Based Cosmic Muon Tracker

Abstract
Primary cosmic rays when interact with our atmosphere, produce a cascade of lighter secondary particles namely pion, kaon, neutrons, muons, electrons, positrons and neutrinos. Muons are one of the most abundant and easily detectable particles at the ground surface using a large variety of particle detectors. Resistive Plate Chambers (RPCs) of in dimension were developed to be used in large scale as the active detector elements in the Iron Calorimeter (ICAL) which was planned to be built by the India-based Neutrino Observatory (INO). As a spin-off of this work, a portable stack of eight, one square foot RPC detectors is developed named as Cosmic Muon Tracker (CMT). It could be used to conduct small-scale particle detector experiments as well as training Students. We will discuss design, integration, characterisation and some of the applications of this detector in this paper.
1 Introduction
When high energy cosmic rays - primarily the protons, hit earth’s atmosphere, they produce showers of secondary particles comprising mostly pions. The pions further decay and produce muons and neutrinos. This shower of particles initially increases and then reduces in number as they travel towards the earth’s surface. As a result, the average flux of muons is about 1 per square centimeter per minute at mean sea level. The muon flux is a function of altitude [1] and their angular momentum depends on the earth’s magnetic field at that location. The wide spectrum of energy and angular distributions makes muons a useful radiation probe for various scientific as well as societal applications such as medical imaging, tomography, structural damage of a bridge/building and various particle physics experiments. The charged cosmic ray muons can be detected using a variety of detectors such as scintillators, Resistive Plate Chambers (RPCs), proportional counters, etc. In general, multiple large area planar detectors arranged in a stacked geometry are used to track muons. RPCs are gaseous detectors [2] which are widely used in particle physics experiments due to their excellent timing and large area coverage, besides ease of fabrication and low cost.
The Iron CALorimeter (ICAL) experiment, proposed by the India-based Neutrino Observatory (INO) collaboration [3], aims to track and study atmospheric neutrinos using a 50-kiloton iron calorimeter. The detector comprises 28,800 Resistive Plate Chambers (RPCs), each measuring , sandwiched between 56 mm thick iron plates to track charged particles produced by neutrino interactions. Each RPC is equipped with a Front-End Data Acquisition module, known as RPC-DAQ, which records the position and timing information of particle trajectories. The collected data is transmitted to back-end servers via an Ethernet interface, with each RPC-DAQ remotely controllable through a dedicated IP address. Event registration and synchronization are managed using a Real-Time Clock system. Based on the detector and instrumentation R&D for the ICAL experiment, a portable detector stack of about one cubic foot in size, capable of tracking cosmic muons in real time is developed. The stack is made up of eight RPCs of about 1 ft × 1 ft in dimension which are filled with a special gas mixture and sealed. Eight Equally segmented signal strips on the top and bottom of the RPC gap of all eight RPCs are readout and processed. Strips on top and bottom layer placed orthogonally and marked as X- and Y-strips. Typically, cosmic ray muons passing through the stack leave tracks in all the eight RPCs and induce signals on the pickup strips of all the planes, which are in the path of the muon trajectory. LEDs - one per strip, mounted on the front-end boards of X- and Y-planes of each RPC in the stack, display the tracks of the cosmic ray muons in real-time as shown in Figure 1. The strip hit information is also stored in an FPGA-based digital Data Acquisition (DAQ) module, which is connected to a computer through Ethernet. The DAQ module also produces the event trigger whenever the input signals match the user-defined trigger criteria. This trigger signal is used to acquire the data from all the strips/channels. The data acquired is transferred to the server and is also displayed as an event track using the on-detector LED matrix. The live display of cosmic muon events can also be ported onto the internet for viewing from anywhere. The low voltages required for powering electronics as well as to generate the high voltage required to bias the RPCs are all locally generated from mains power, thus making this tracker highly portable. The paper describes design and fabrication of the detectors, electronics, trigger and data acquisition systems, performance of the stack as well as its other potential applications.
2 Detector Geometry
As shown in Figure 2, the Cosmic Muon Tracker (CMT) consists of 8 RPCs stacked vertically and each separated by . Each layer is supported by thick aluminum honeycomb panels with C-type aluminum channels providing vertical support. Each RPC is centered in its layer using small L-shaped aluminum brackets. The overall size of the stack is with each RPC detector measuring Out of area of the RPC, only is the active detector surface for tracking muons. The X- and Y-sides of the stack are dedicated to the front-end modules and LED displays while the back of the X-side is used for cabling the high voltage distribution. A manifold made of PVC pipes used for gas distribution is mounted on the rear of the Y-side. The top surface of the stack is used for accommodating the data acquisition (DAQ) module along with low and high voltage modules. High-density differential signal cables connect each front-end module to the DAQ module and the low-voltage supply for each front-end is routed through the empty slots of the C-type channels from top to bottom of the stack. Transparent acrylic sheets are used to cover all sides of the stack to protect the front-end modules and RPCs.
3 Working principle of the RPC detector
RPCs consist of two thick parallel glass plates with a bulk resistivity of about . A semi-conductive paint is coated on one side of the glass plates. These plates are mounted parallel to each other - with uncoated sides facing each other, with a gap of between them which was maintained using poly-carbonate button-shaped spacers. The volume thus formed is sealed using 3M DP-190 glue on all four sides. This assembly is called the RPC gas gap. A bias voltage of about is applied across the gas gap using HV contacts pasted onto the conductive coating surfaces. Signal pickup is achieved using copper strip-based honeycomb panels placed on either side of the gas gap in an orthogonal configuration. Thin Mylar sheets are used to isolate the pickup strips from the conductive coating carrying high voltage.
The RPC gas gap is fitted with gas inlet and outlets to flow the ionization gas mixture consisting of R-134a (95%), isobutane(4.2%) and SF6(0.3%). through the chamber. Main components of an RPC detector are shown in Figure 3. RPC operates on the principle of ionization as shown in the Figure 4. A uniform electric field is created between the two plates inside the gas volume by applying a bias voltage across the chamber. When a charged particle (such as a cosmic ray muon) passes through the chamber, it ionizes the gas and thus generates free electrons and positive ions. The electrons are accelerated towards the anode due to the electric field, triggering a series of secondary ionizations on the way and thus leading to an avalanche.
This avalanche induces an electric signal on the copper strips attached to the external pickup panels. Each side of the RPC is divided into eight copper strips, each of width and a pitch of . The resistive nature of the glass plates ensures that the discharge is localized, and prevents it from spreading across the entire plate. The pickup strips are appropriately terminated to readout the signal from one end only. The signals are then transmitted through coaxial cables to the front-end electronics boards for further processing.
4 RPC Production and Testing
Before assembling the cosmic muon tracker, each RPC was fabricated and quality checked [4] with assistance from summer project students who were invited from various institutions across India. The students were assigned tasks such as measuring the surface resistivity of the conductive coating, assembling the RPC, leak rate estimation of the gas gaps, performing V-I characteristics and final characterization of the detectors.
4.1 Surface resistivity measurement
In order to ensure uniform distribution of the bias voltage and thus electric field across the chamber, the conductive coating must have uniform surface resistivity [5]. This is measured using a specially developed contact jig and a multi-meter. The jig consists of two copper coated brass pieces separated by a plastic spacer of the same length as that of the brass pieces. The enclosed paint surface by the jig thus forms a square. As shown in Figure 5 the jig is placed in 2 different orientations – called “H” and “I” in a matrix on the painted surface.
Figure 6 shows the surface resistivity measurements for an RPC in both orientations on top and bottom surfaces. Similar measurements were recorded for all the eight RPCs. Most of the painted glasses showed a uniform surface resistivity of about . It may be noted that the quantity measured by the multi-meter is resistance, while it actually represents a quantity called “resistance per square” due to construction of the jig as described above.
4.2 Leak test of RPC Gaps
The fabricated gas gaps must be gas-tight else the performance of the RPC could be compromised due to the contaminated gas mixture. To ensure that the RPCs are leak free, leak test measurements [6] were conducted using an Arduino-based system. The outlet of the RPC was connected to a pressure sensor linked to a NodeMCU microcontroller Wi-Fi module which transmits the pressure readings inside the chamber directly to a server. The inlet of the RPC was connected to a gas supply with a needle valve to finely adjust the pressure filling the chamber with nitrogen gas.
Another NodeMCU based pressure sensor module was used to measure the atmospheric pressure readings and transmitted to the server. A Python script was used to store the data and display the pressure difference over time. The pressure inside the RPC was increased above the atmospheric pressure and tightly sealed. The pressure inside the RPC and pressure difference are continuously monitored and logged at least for 5 hours. If the pressure remained stable within its initial range, then the chamber was deemed to be leak-free and is ready for further fabrication. The temperature inside the chamber was also plotted as a function of time.
A temperature variation of 22.5∘C. to 24.5∘C was observed due to the ON-OFF duty cycle of the lab air conditioners. In case a gas leak was found, a hand-held gas leak tester was used to locate the exact position of the leak and the RPC was re-glued. Typical leak test plots include temperature variation (mostly due to air conditioning), atmospheric pressure (black), pressure inside the chamber (red), and the pressure difference as shown in the bottom plot of Figure 7. The leak-test data similar to that of the electrode paint resistance measurements was documented for all RPCs for future reference.
4.3 I-V Characterization
Figure 8 shows the equivalent circuit of an RPC gas gap. At lower voltages, the current flows from the top plate to the bottom plate through the spacers with the resistance denoted as . When the RPC attains the operating potential, an electric field is established between the top and bottom plates and the current begins to flow through the gas medium whose resistance is represented as .
For the I-V characteristics test, the RPCs were connected to a differential high voltage supply and the voltage was gradually increased from to in steps of for each polarity. Channel A corresponds to positive polarity and channel B to negative polarity, meaning that a total potential difference of is applied when both channels reach a voltage of each. Current readings were recorded at each step after allowing the currents to stabilize. Figure 9 shows the I-V characteristics plots of both polarities of all the 8 RPCs used in the CMT.
4.4 RPC Detector Characterization
Efficiency of the RPCs are measured using a set of reference scintillator paddles which are optically coupled to a photomultiplier tube (PMT). The scintillator paddles receive high voltage from the HV distribution box while the RPC is supplied with bias supply separately. The signal outputs from the scintillator paddles are fed to a discriminator module where adjustments to the threshold and pulse width can be made. Since RPC pulses are small - typically in the millivolt range, they are amplified using a Hybrid Micro Circuits (HMC) preamplifier, a transistor-based voltage amplifier with a gain of about 80 before feeding them into the discriminator modules.
Three scintillator trigger paddles are aligned on top and bottom of the RPC covering a specific area of the RPC strip. The overall test setup is shown in Figure 10. The successful muon events generate signals in all the three paddles and RPC strips. The efficiency of the RPC is determined by the ratio of the 4Fold (P1 AND P2 AND P3 AND RPC strip) to the 3Fold (P1 AND P2 AND P3). The HV bias of the RPC is slowly increased in steps of varying from to . At each SET voltage, the counting rates of the RPC strip signal, 3Fold and 4Fold are measured using a visual pre-scaler. The efficiency of an RPC is plotted as shown in Figure 11.
Figure 12 shows the noise rate of one of the RPC channels. It is essential to ensure that the noise rate of the RPC is minimal before conducting this measurement. Noise rate is one of the basic health monitoring parameters of the RPC. Noise rate of an RPC must be stable for proper tracking of muons through the RPC stack.
5 Stacking of CMT
After characterizing all eight RPC detectors, they were prepared for assembly into the CMT stack. Each RPC was placed on an aluminium honeycomb tray for rigidity and reduced weight. The detectors were fitted to the base plate using custom-fabricated “Z” clamps. Each of the eight layers was then mounted sequentially onto a rigid aluminium frame made from extruded sections as shown in Figure 13. Power cables for all the front-end (FE) boards were routed through the slot in the extruded section. Aluminium mounts for the FE boards on the X- and Y-sides were also fabricated to ensure rigid mounting. The complete stack of RPC layers was equipped with all necessary electronic components including FE boards, power supply boards, RPC data acquisition systems and gas supply manifolds. Cable harness and routing were done with all components connected to a common ground point using wires and copper strips to ensure a reliable ground reference. Signal cables were properly shielded to support high speed signal transmission and reduced noise.
Gas inlet of each RPC is connected in cascade to support flow of gas in all the RPCs uniformly. The pressure inside the RPC chambers were stable over time due to tight sealing of the gas inlets and outlets. The high voltage required for operation of the RPC is generated using an onboard power supply, thus enabling the CMT to be operated directly from an A.C. mains supply at any location.
6 Trigger and Data Acquisition System
To ensure the portability of the tracker, all the electronics comprising of Data Acquisition Module (DAQ), 16 Front-end circuits (FE), Low Voltage-High voltage (LVHV) module, power lines, signal cables and gas manifolds are harnessed carefully within the tracker geometry. The DAQ, LVHV and power supply modules are placed on top of the stack as shown in Figure 14.
The FE Boards on X- and Y-side are well interfaced with the FPGA-based Data Acquisition module named RPC-DAQ [7] shown in the Figure 15, which was originally developed as a digital Front-End for RPCs for the ICAL experiment. RPC-DAQ is capable of handling 128 inputs covering 64 channels from X-side and 64 channels from Y-side.
When a charged particle traverses the CMT, the RPCs strips in each layer and which are in the path of muon trajectory produce induced signals in the X- and Y-sides. The FE modules shown in Figure 16, in each layer amplify these signals and discriminate them above a fixed threshold. The resultant LVDS (Low Voltage Differential Signaling) logic signals are then sent to the RPC-DAQ module placed on top of the CMT unit.
These signals are processed within a few nanoseconds and the muon track is identified by a coincidence hardware logic shown in Figure 17. On successful coincidence – indicating passage of a cosmic muon through the stack, a trigger signal is generated to latch the event track hit information responsible for the trigger. This latched event is then sent back to the FEs and LEDs of corresponding RPC strips are lit, thus indicating live the cosmic track. This way, every muon generated track is displayed by the corresponding LEDs. Until the next trigger forms, the previous latched state is maintained in the LEDs of all layers.
Also, all the incoming signals to the FPGA are sent to a High-Performance Time to Digital Converter (HPTDC) [8] for time-stamping the input signals with resolution. This helps in tracking muons with accurate directional measurement. Also, a Real-Time Clock (RTC) is used to timestamp every muon event. Apart from LED tracking, the DAQ FPGA acquires position, timing and RTC information for every event and transfers the same to the computer using standard Ethernet protocols. Necessary hardware logics are implemented to continuously monitor the health of the detector. Both Event and health monitoring data are acquired by a "C" language-based Back-End (BE) server software. The BE server displays some basic information to the user from time to time and stores the acquired data. The BE server is equipped with features to configure various DAQ settings like trigger layers, RTC and channel masking at the start of the run. These features enhance the functionality of the user-friendly data acquisition for the tracker.
7 System Efficiency, Technical Challenges and Student Training
The Cosmic Muon Tracker (CMT) was evaluated extensively to assess detector performance, data acquisition stability, portability and its utility in student training. The primary metric of interest was the efficiency of the Resistive Plate Chambers (RPCs), determined from muon track reconstruction. Muon tracks were fitted using a least-square method to estimate expected hit positions across each RPC layer. The measured hits were then compared against the expected trajectory to compute efficiency. Most layers demonstrated high efficiency, with six out of eight RPCs exceeding (90%) as summarized in Table 1. Instances of reduced efficiency during field deployment prompted RPC repair or replacement to maintain system reliability.
| Layer0 | Layer1 | Layer2 | Layer3 | Layer4 | Layer5 | Layer6 | Layer7 |
| 96.6 | 90 | 48.3 | 90 | 93.2 | 82.5 | 96.5 | 70.6 |
The LED event display integrated into the Data Acquisition (DAQ) logic provides real-time visual feedback by illuminating LEDs corresponding to muon hits Figure 18. Its stability under varying environmental conditions such as temperature, humidity and pressure was closely monitored. These parameters influence the RPC noise rate and signal multiplicity, thereby impacting the clarity of track visualization. Continuous logging of strip noise rates allowed the detection of abnormalities and ensured consistent detector performance during long-term operation.
The CMT system was also deployed as a platform for student engagement in hands-on detector development and outreach as shown in Figure 19. Graduate and undergraduate students were trained in operating the DAQ system, troubleshooting noise issues, operating RPC voltages, and analyzing muon data. This practical exposure to high-energy physics instrumentation significantly contributed to technical skill development.
CMT serves as a scalable and portable platform for RPC detector development and student-led particle physics research as well as used in INO-ICAL project outreach programs such as Vigyan Samagam.
8 Conclusion
CMT is a gaseous detector that requires regular monitoring to ensure optimal performance. The tracker is flushed with fresh gas mixture whenever its performance degrades. Its performance is also tested under various ambient conditions. The compact and portable design of the CMT makes it ideal for measuring muon flux at different altitudes. With a timing accuracy of , it can determine the direction of muon paths and thus can be used for tomography. It can serve as a simple counter for background measurements in mines, etc. The CMT can also function in both vertical and horizontal orientations allowing for comparisons of horizontal and vertical fluxes at different locations as well as any zenith angle.
Acknowledgments
We extend our gratitude to Antara Prakash and Pratik Barve, intern students, for their valuable contributions to the initial development of the detector and electronics. We also sincerely thank the participants of the RPC lab from the SERB School 2019 held at TIFR for their support in fabricating the RPC detector. Also, we sincerely thank all our present and former INO colleagues especially Puneet Kanwar Kaur, Suraj Kole, Sagar Sonavane, Rajkumar Bharathi for technical support during the design, also we would like to thank members of TIFR, namely K.C. Ravindran, Darshana Gonji, Santosh Chavan and Vishal Asgolkar who supported testing and commissioning. We would like to thank Former INO students Neha Panchal, Abhijit Garai for extending their support during the development. Also, we like to thank former INO directors V.M Datar abd N.K. Mondal for his continuous encouragement and guidance.